Check whether the function, ,f defined below, is uniformly continuous or not:



f(x)=x^(1/2), x∈[1,2]

1
Expert's answer
2022-03-15T13:11:09-0400

ANSWER: The function "f(x)=x^ \frac{1}{2}" is uniformly continuous on [1,2]

EXPLANATION 1.

Let "\epsilon>0" . We need to show that exists "\delta>0" such that

"x,y \in[1,2]" and "\left | x-y \right |<\delta" imply "|f(x)-f(y)|<\epsilon" .

We have

"f(x)-f(y) =\sqrt{x}-\sqrt{y}=\frac{\left ( \sqrt{x} -\sqrt{y}\right )\cdot\left ( \sqrt{x}+\sqrt{y} \right )}{\sqrt{x}+\sqrt{y}}=\frac{x-y}{\sqrt{x}+\sqrt{y}}" .

Because "\sqrt{x}\geq1, \sqrt{y}\geq1" , then "\frac{\left | x-y \right |}{\sqrt{x}+\sqrt{y}}\leq\frac{\left | x-y \right |}{2}" . If "\delta=2\epsilon" and "|x-y| \leq\delta" , then "\left | f(x)-f(y) \right |\leq\frac{\left | x-y \right |}{2}\leq\frac{\delta}{2}=\frac{2\epsilon}{2}=\epsilon" . So, we have shown, by the definition, that "f" is uniformly continuous on [1,2].

EXPLANATION 2.

The function "f" is continuous on a closed interval [1,2] , so by Cantor's theorem the function "f" is uniformly continuous on [1,2].